The time spent waiting in the line is approximately normally distributed. The mean waiting time is 6 minutes and the variance of the waiting time is 4. Find the probability that a person will wait for more than 7 minutes. Round your answer to four decimal places.
Answer #1

Solution :

Given ,

mean = = 6

2 =4

standard deviation = = 2

P(x >7 ) = 1 - P(x< 7)

= 1 - P[(x -) / < ()7-6 / 2]

= 1 - P(z <0.5)

Using z table

= 1 - 0.6915

= 0.3085

probability= 0.3085